OpenStudy (fanduekisses):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11)

OpenStudy (anonymous):

foci c =\(\pm11\) vertices (0, +/- 2) --> hyperbola forms the form of \(\dfrac{y^2}{a^2}-\dfrac{x^2}{b^2}=1\) you have a = 2, find b by c^2= a^2+ b^2 and plug them in

OpenStudy (fanduekisses):

so it is \[\frac{ y ^{2}}{ 4 } - \frac{ x^{2} }{121}=1\]